From GlueXWiki
Jump to: navigation, search

August 30, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
  2. Cathode corrosion
    • Recent test results: [1]
  3. Conductive tape studies: [2], [3], [4],[5], [6] (Vlad)
  4. Engineering (Bill)
    • Cathode strong-backs
  5. Electronics (Chris, Nick)
    • PCBs
  6. Third package tests FDC E-log, Entry 71(Beni)
  7. TDR, pages 132-145 (Lubomir)
  8. Other


Participants: Fernando, Bill, Dave, Elton, Simon, Chris, Vlad, Beni, and Lubomir


- Dave: working on all the cathodes for package #1. Using the new tool to hold the cathodes flat, two cathodes moved to tables and the cards are being glued there. Will have cathodes for the first, second, and third cells by next Tuesday, next Thursday, and next next Tuesday respectively. This is assuming one cathode per day, which is limited by the tools that we use now.

- Wire frames: Chris will put the resistors on the first PCB set next week. Then we will start laminating the wire plane #29, we have a frame for it. The other three G10 wire frames will come from the machine shop in ~2 weeks. Fernando and Bill confirmed that the new PCBs can be used.

- Roger got quotations from Allflex and Flexible Circuits, the two Minnesota companies that made the cathode foils and the rigid-flex assemblies. According to Allflex they have material for 18-19 foils and we asked for 20. We will need the foils for some of the cathodes on package #2. Based on Dave's estimations for the refurbishment of the first package and then completion of the fourth one, we will need the foils sometime in November-December. In any case it's better to submit the PRs now and contact Eugene to use funds (~$26K) from FY12, if possible?

Cathode corrosion

- The last test with the vacuum chamber: we had a sample with Viton+Apiezon, three O-ring pieces across a strip using the strong (red) clamp and keeping it for one week at high temperature and humidity, the longest time so far. The picture of the sample is linked above: copper changed color everywhere except at the place where the Apiezon was applied. Several dark lines are visible at the O-ring positions. Two on the sides are darker, probably they took the main pressure from the clamp. Only in the middle we see some very small holes through the copper, visible only with a good back-light. Olga will do SEM analyses.

- This is the first time we had dark traces under the Vition+Apiezon O-rings, though the conditions of the above test were extreme and we see corrosion/oxidation everywhere on the sample. To make sure we have the right pressure on the Viton we started a new test with a real spacer ring "inserted" through the vacuum chamber and applying pressure on it with a lead brick.

Conductive tape studies

- Vlad did the same ball counting studies as before, but with another cathode (#4, one of the worst: top of former package #1) this time using improved technique in removing the conductive tape and measuring the resistances. The correlation between the resistances and the ball numbers is much more evident now (profile plot above). Again we have about 0.5-2 Ohms per ball and all the resistances above several Ohms (~10-15%) can be explained by loose contacts between the balls and the cathode traces. The highest resistance on this cathode was 50 Ohms.


- The cathode strong-back tool needs more suction cups. The six that we have now are enough for type-1 but not for type-2/3. Will have them on Tuesday and then can continue with the cathode production. Bill suggested installing a gauge to make sure the vacuum is OK before lifting the cathode.


- Fernando did some organization of the files for all the FDC electronic parts; he also got the latest versions from Roger. Now the production files are separated from the older versions.

- Fernando explained the status with the fADC125. Beni and Cody did some tests: the problem is the bus communication, not in between the FPGAs.

Third package tests

- Third package was moved to 126 last Friday, started bubbling immediately, flushed with gas (40/60 Ar/CO2) over the weekend.

- Beni: Nominal HV (+2100/-500V) was applied starting Monday without problems, except one negative (field wires) HV sector on the bottom (#6) cell that trips from time to time. For the overnight measurements there we applied +2250V on the sense wires only.

- Linux DAQ works now for the TDCs only, there was a problem with byte swapping. First results are linked above from the FDC E-log:

  1. Efficiency: much better than with the first package (means oxygen % very low) in the 95-100% range, and if needed we can increase the HV. However more studies are needed; the efficiency is not so great when applying timing cuts.
  2. Noise in the second cell. After the meeting it turned out that we missed one ground connection there. After installing it the noise seamed to disappeared.
  3. TDC spectra look nice and uniform with two peaks well defined, corresponding to the position of the sense and field wires. The spectrum for the HV sector without field voltage is much wider.
  4. Missing half of cell #3 due to translator board problem, now being repaired by Cody.


- Latest version posted. FDC electronics subsection: we should give details up to the pre-amp level. The rest of the electronics (fADC125, F1TDC, ...) should be in the "Electronics" section.